# Parallel transport on matrix manifolds and Exponential Action

### v1

@@ 3. Transposable spaces, trace and Frobenius forms, and a family of tractable metrics
(23) $\langle\mathrm{P}_{\mathtt{a}}\mathtt{b},\mathtt{c}\rangle_{\vec{\beta}}+\langle\mathrm{P}_{\mathtt{a}}\mathtt{c},\mathtt{b}\rangle_{\vec{\beta}}=0.$
It is known $\mathtt{T}(\gamma)_{0}^{t}$ is an isometry, which also follows from eq. 23.
@@
Set $\mathcal{I}_{\vec{\beta}}:\omega\mapsto\beta_{0}(\omega-\omega_{\mathfrak{a}})^{\mathsf{T}}-\beta_{1}\omega_{\mathfrak{a}}^{\mathsf{T}}$, then $\langle\omega_{1},\omega_{2}\rangle_{\vec{\beta}}=\langle\omega_{1},\mathcal{I}_{\vec{\beta}}\omega_{2}\rangle_{\mathcal{F}}$. As an operator on $\mathcal{E}=\mathbb{R}^{n\times n}$, the operator $\mathcal{I}_{\vec{\beta}}\circ\mathsf{T}$ has two eigenvalues, $\beta_{0}$ with eigenspace $(\operatorname{I}_{\mathcal{E}}-\mathrm{p}_{\mathfrak{a}})\mathcal{E}$, (recall $\mathrm{p}_{\mathfrak{a}}$ is the projection to $\mathfrak{a}$) and $-\beta_{1}$ with eigenspace $\mathfrak{a}$, both are nonzero by assumption. Thus, the operator $\mathcal{I}_{\vec{\beta}}$ and the form $\langle\rangle_{\vec{\beta}}$ are nondegenerate.
For eq. 19, decompose $\mathtt{a}=\mathtt{a}_{\mathfrak{a}}+\mathtt{a}_{\mathfrak{a}_{{}_{\perp}}},\mathtt{b}=\mathtt{b}_{\mathfrak{a}}+\mathtt{b}_{\mathfrak{a}_{{}_{\perp}}}$, and substitute in, note $[\mathtt{a}_{\mathfrak{a}},\mathtt{b}_{\mathfrak{a}}]+[\mathtt{b}_{\mathfrak{a}},\mathtt{a}_{\mathfrak{a}}]=0$, the remaining part is in $\mathfrak{a}_{{}_{\Join}}$ by eq. 13.
To show eq. 18 defines a connection, we need to show $X^{-1}(\nabla_{\mathtt{Y}}\mathtt{Z})_{X}\in\mathfrak{g}$, and we only need to show the following expression (dropping the variable X in the functions for brevity) is in $\mathfrak{g}$ when evaluated at $X\in\mathrm{G}$, as the remaining terms are already in $\mathfrak{g}$
$X^{-1}\mathrm{D}_{\mathtt{Y}}\mathtt{Z}-\frac{1}{2}(X^{-1}\mathtt{Y}X^{-1}\mathtt{Z}+X^{-1}\mathtt{Z}X^{-1}\mathtt{Y})\in\mathfrak{g}.$
Here, set $\mathtt{y}(X)=X^{-1}\mathtt{Y}(X),\mathtt{z}(X)=X^{-1}\mathtt{Z}(X)$, then $\mathtt{y}$ and $\mathtt{z}$ are two $\mathfrak{g}$-valued functions on $\mathrm{G}$, the expression becomes
(24) $X^{-1}(X\mathtt{y}\mathtt{z}+X\mathrm{D}_{\mathtt{Y}}\mathtt{z}-\frac{1}{2}(\mathtt{y}\mathtt{z}+\mathtt{z}\mathtt{y}))=\mathrm{D}_{\mathtt{Y}}\mathtt{z}+\frac{1}{2}[\mathtt{y},\mathtt{z}]\in\mathfrak{g}$
where $\mathrm{D}_{\mathtt{Y}}\mathtt{z}\in\mathfrak{g}$ because of linearity, while $[\mathtt{y},\mathtt{z}]\in\mathfrak{g}$ because $\mathfrak{g}$ is closed under the Lie bracket. As $\Gamma$ is symmetric in $\mathtt{Y}$ and $\mathtt{Z}$, it remains to check for metric compatibility, or to show $C:=\mathrm{D}_{\mathtt{Y}}\langle\mathtt{Z},\mathtt{Z}\rangle_{\vec{\beta}}-2\langle\mathtt{Z},\nabla_{\mathtt{Y}}\mathtt{Z}\rangle_{\vec{\beta}}$ vanishes assuming eq. 18. Expand, we have
$\begin{gathered}C=\beta_{0}\mathrm{D}_{\mathtt{Y}}(\operatorname{Tr}\mathtt{z}^{2})-(\beta_{0}+\beta_{1})\mathrm{D}_{\mathtt{Y}}(\operatorname{Tr}\mathtt{z}_{\mathfrak{a}}^{2})\\ -2\beta_{0}\operatorname{Tr}\mathtt{z}X^{-1}\nabla_{\mathtt{Y}}\mathtt{Z}+2(\beta_{0}+\beta_{1})\operatorname{Tr}\mathtt{z}_{\mathfrak{a}}(X^{-1}\nabla_{\mathtt{Y}}\mathtt{Z})_{\mathfrak{a}}.\end{gathered}$
Using eq. 24 to simplify $X^{-1}\nabla_{\mathtt{Y}}\mathtt{Z}$, the second line becomes
$\begin{gathered}-\beta_{0}\left(2\operatorname{Tr}\mathtt{z}\mathrm{D}_{\mathtt{Y}}\mathtt{z}+\operatorname{Tr}\mathtt{z}[\mathtt{y},\mathtt{z}]+(1+\beta)\operatorname{Tr}\mathtt{z}\left([\mathtt{y}_{\mathfrak{a}},\mathtt{z}]+[\mathtt{z}_{\mathfrak{a}},\mathtt{y}]\right)\right)\\ +(\beta_{0}+\beta_{1})\left(2\operatorname{Tr}\mathtt{z}_{\mathtt{a}}\mathrm{D}_{\mathtt{Y}}\mathtt{z}_{\mathtt{a}}+\operatorname{Tr}\mathtt{z}_{\mathtt{a}}[\mathtt{y},\mathtt{z}]_{\mathtt{a}}+(1+\beta)\operatorname{Tr}\mathtt{z}_{\mathtt{a}}\left([\mathtt{y}_{\mathfrak{a}},\mathtt{z}]+[\mathtt{z}_{\mathfrak{a}},\mathtt{y}]\right)\right)_{\mathtt{a}}.\end{gathered}$
The last term vanishes by eq. 19. Putting together and cancelling terms with $\mathrm{D}_{\mathtt{Y}}\mathtt{z}$
$\begin{gathered}C=-\beta_{0}\left(\operatorname{Tr}\mathtt{z}[\mathtt{y},\mathtt{z}]+(1+\beta)\operatorname{Tr}\mathtt{z}\left([\mathtt{y}_{\mathfrak{a}},\mathtt{z}]+[\mathtt{z}_{\mathfrak{a}},\mathtt{y}]\right)\right)\\ +(\beta_{0}+\beta_{1})\left(\operatorname{Tr}\mathtt{z}_{\mathtt{a}}[\mathtt{y},\mathtt{z}]_{\mathtt{a}}\right)\\ =-(\beta_{0}+\beta_{1})\operatorname{Tr}\mathtt{z}[\mathtt{z}_{\mathfrak{a}},\mathtt{y}]+(\beta_{0}+\beta_{1})\left(\operatorname{Tr}\mathtt{z}_{\mathtt{a}}[\mathtt{y},\mathtt{z}]_{\mathtt{a}}\right)\\ =-(\beta_{0}+\beta_{1})\left(\operatorname{Tr}\mathtt{z}[\mathtt{z}_{\mathfrak{a}},\mathtt{y}]-\operatorname{Tr}\mathtt{z}_{\mathtt{a}}[\mathtt{y},\mathtt{z}]\right)=0,\end{gathered}$
where we note $\operatorname{Tr}\mathtt{z}[\mathtt{y},\mathtt{z}]=0=\operatorname{Tr}\mathtt{z}[\mathtt{y}_{\mathfrak{a}},\mathtt{z}]$ by eq. 10, $\operatorname{Tr}\mathtt{z}_{\mathtt{a}}[\mathtt{y},\mathtt{z}]=\operatorname{Tr}\mathtt{z}_{\mathtt{a}}[\mathtt{y},\mathtt{z}]_{\mathfrak{a}}$ since $\mathrm{p}_{\mathfrak{a}}$ is a projection, then finish the cancelation by eq. 10 again.
To verify that eq. 20 satisfies the geodesic equation, with $\mathtt{a}=X^{-1}\xi$, we have
(25) $\dot{\gamma}(t)=X\operatorname{expm}(t\left(-(1+\beta)\mathtt{a}_{\mathfrak{a}}+\mathtt{a}\right))\mathtt{a}\operatorname{expm}(t(1+\beta)\mathtt{a}_{\mathfrak{a}}),$
(26) $\ddot{\gamma}(t)=X\operatorname{expm}(t\left(-(1+\beta)\mathtt{a}_{\mathfrak{a}}+\mathtt{a}\right))\left(\mathtt{a}^{2}+(1+\beta)[\mathtt{a},\mathtt{a}_{\mathfrak{a}}]\right)\operatorname{expm}(t(1+\beta)\mathtt{a}_{\mathfrak{a}}).$
From here $\gamma^{-1}\dot{\gamma}=U^{-1}\mathtt{a}U$ for $U=\operatorname{expm}(t(1+\beta)\mathtt{a}_{\mathfrak{a}})$, while
$D:=\gamma(t)^{-1}\Gamma(\dot{\gamma}(t),\dot{\gamma}(t))_{X=\gamma(t)}=-U^{-1}\mathtt{a}^{2}U+(1+\beta)U^{-1}[\mathtt{a}_{\mathfrak{a}},\mathtt{a}]U)$
since from the expression for $U$, $U^{-1}\mathtt{a}_{\mathfrak{a}}U=\mathtt{a}_{\mathfrak{a}}$, thus, $D$ simplies to the above. This shows $\ddot{\gamma}(t)+\Gamma(\dot{\gamma},\dot{\gamma})=0$, hence $\gamma(t)$ is the geodesic with $\gamma(0)=X,\dot{\gamma}(0)=X\mathtt{a}=\xi$.
For parallel transport, assume the vector field $\Delta(t)$ over the curve $\gamma(t)$ satisfies the transport equation $\dot{\Delta}+\Gamma(\dot{\gamma},\Delta)=0.$ Let $W(t)=U(t)\gamma(t)^{-1}\Delta(t)U(t)^{-1}$, where $U(t)=\operatorname{expm}(t(1+\beta)\mathtt{a}_{\mathfrak{a}})$. We will show $W(t)$ satisfies a linear differential equation with constant coefficients. Express $\Delta$ in terms of $W(t)$ and take time derivative
$\begin{gathered}\Delta=\gamma U^{-1}WU,\\ \dot{\Delta}=\dot{\gamma}U^{-1}WU+(1+\beta)\gamma U^{-1}[W,\mathtt{a}_{\mathfrak{a}}]U+\gamma U^{-1}\dot{W}U\\ =\gamma U^{-1}(\mathtt{a}W+(1+\beta)[W,\mathtt{a}_{\mathfrak{a}}]+\dot{W})U,\end{gathered}$
since $\dot{\gamma}=\gamma U^{-1}\mathtt{a}U$, thus, $\dot{\gamma}U^{-1}WU=\gamma U^{-1}\mathtt{a}UU^{-1}WU=\gamma U^{-1}\mathtt{a}WU$. Next,
$\Gamma(\dot{\gamma},\Delta)=-\frac{1}{2}\gamma U^{-1}(\mathtt{a}W+W\mathtt{a})U+\frac{1+\beta}{2}\gamma U^{-1}\left([\mathtt{a}_{\mathfrak{a}},W]+[W_{\mathfrak{a}},\mathtt{a}]\right)U.$
From here, $U\gamma^{-1}(\dot{\Delta}+\Gamma(\dot{\gamma},\Delta))U^{-1}=0$ gives the equation for $W$
$\dot{W}+\frac{1}{2}[\mathtt{a},W]+\frac{1+\beta}{2}\left([W_{\mathfrak{a}},\mathtt{a}]+[W,\mathtt{a}_{\mathfrak{a}}]\right)=0.$
Thus, $W(t)=\operatorname{expv}(t\mathrm{P}_{\mathtt{a}},W_{0})$ with $W_{0}=X^{-1}\eta$, implying eq. 21.
To prove eq. 22, we use the equality $\operatorname{Tr}\mathtt{b}_{1}^{\mathsf{T}}[\mathtt{b}_{2},\mathtt{g}]=\operatorname{Tr}\mathtt{b}_{2}^{\mathsf{T}}[\mathtt{b}_{1},\mathtt{g}^{\mathsf{T}}]$ for $\mathtt{b}_{1},\mathtt{b}_{2},\mathtt{g}\in\mathbb{R}^{n\times n}$ by taking transpose then use eq. 10. Thus, for $\mathtt{c}\in\mathcal{E}=\mathbb{R}^{n\times n}$,
$\begin{gathered}\operatorname{Tr}\mathtt{c}^{\mathsf{T}}\left(\frac{1}{2}[\mathtt{b},\mathtt{a}]+\frac{1+\beta}{2}\left(-[\mathtt{b}_{\mathfrak{a}},\mathtt{a}]-[\mathtt{b},\mathtt{a}_{\mathfrak{a}}]\right)\right)\\ =\frac{1}{2}\operatorname{Tr}\mathtt{b}^{\mathsf{T}}[\mathtt{c},\mathtt{a}^{\mathsf{T}}]-\frac{1+\beta}{2}\left(\operatorname{Tr}\mathtt{b}_{\mathfrak{a}}^{\mathsf{T}}[\mathtt{c},\mathtt{a}^{\mathsf{T}}]+\operatorname{Tr}\mathtt{b}^{\mathsf{T}}[\mathtt{c},\mathtt{a}_{\mathfrak{a}}^{\mathsf{T}}]\right).\end{gathered}$
Using $\operatorname{Tr}\mathtt{b}_{\mathfrak{a}}^{\mathsf{T}}[\mathtt{c},\mathtt{a}^{\mathsf{T}}]=\operatorname{Tr}\mathtt{b}_{\mathfrak{a}}^{\mathsf{T}}[\mathtt{c},\mathtt{a}^{\mathsf{T}}]_{\mathfrak{a}}=\operatorname{Tr}\mathtt{b}^{\mathsf{T}}[\mathtt{c},\mathtt{a}^{\mathsf{T}}]_{\mathfrak{a}}$, we get the expression for $\mathrm{P}^{\mathsf{T}}_{\mathtt{a}}$.
We prove eq. 23 by a direct calculation, note $\operatorname{Tr}\mathtt{g}_{\mathfrak{a}}\mathtt{c}_{\mathfrak{a}}=\operatorname{Tr}\mathtt{g}\mathtt{c}_{\mathfrak{a}}$ for $\mathtt{g}\in\mathfrak{g}$

### v2

@@ 4.2. Deformation using an intertwining operator on a subalgebra
(21) $\langle(\mathtt{X}_{1},\hat{\mathcal{B}}\mathcal{P}_{\mathfrak{a}}\mathtt{X}_{1}),(\mathtt{X}_{2},\hat{\mathcal{B}}\mathcal{P}_{\mathfrak{a}}\mathtt{X}_{2})\rangle_{\tilde{\mathfrak{m}}}:=\beta_{0}\langle\mathtt{X}_{1},\mathtt{X}_{2}\rangle-\langle\beta_{0}^{2}\mathcal{B}^{-1}(\operatorname{I}_{\mathfrak{a}}+\beta_{0}^{-1}\mathcal{B})\mathcal{P}_{\mathfrak{a}}\mathtt{X}_{1},\mathcal{P}_{\mathfrak{a}}\mathtt{X}_{2}\rangle$
satisfies the naturally reductive condition, inducing the metric eq. 18 on $\mathcal{G}$.
If $\operatorname{I}_{\mathfrak{a}}+\beta_{0}^{-1}\mathcal{B}$ is invertible then for $(\mathtt{X}_{1},\mathtt{A}_{1}),(\mathtt{X}_{2},\mathtt{A}_{2})\in\mathfrak{g}\times\mathfrak{a}$, define the metric
(22) $\langle(\mathtt{X}_{1},\mathtt{A}_{1}),(\mathtt{X}_{2},\mathtt{A}_{2})\rangle_{\tilde{\mathcal{G}},\vec{\beta}}:=\beta_{0}\langle\mathtt{X}_{1},\mathtt{X}_{2}\rangle-\langle\beta_{0}\mathcal{B}(\mathcal{B}+\beta_{0}\operatorname{I}_{\mathfrak{a}})^{-1}\mathtt{A}_{1},\mathtt{A}_{2}\rangle.$
It is bi-invariant on $\tilde{\mathcal{G}}$ and restricts to eq. 21. The mapping $\mathfrak{q}:\tilde{\mathcal{G}}\to\mathcal{G}$ defined by $\mathfrak{q}:(X,A)\mapsto XA^{-1}\in\mathcal{G}$ is a pseudo-Riemannian submersion between $\langle,\rangle_{\tilde{\mathcal{G}},\vec{\beta}}$ and $\langle,\rangle_{\vec{\beta}}$.
Remark 1.
1. Since $\mathcal{B}$ is symmetric and $\mathcal{P}_{\mathfrak{a}}$ is a projection, in eq. 18
(23) $\langle(\mathcal{B}+\beta_{0}\operatorname{I}_{\mathfrak{a}})\mathcal{P}_{\mathfrak{a}}\mathtt{X}_{1},\mathcal{P}_{\mathfrak{a}}\mathtt{X}_{2}\rangle=\langle\mathcal{P}_{\mathfrak{a}}\mathtt{X}_{1},(\mathcal{B}+\beta_{0}\operatorname{I}_{\mathfrak{a}})\mathcal{P}_{\mathfrak{a}}\mathtt{X}_{2}\rangle=\langle\mathtt{X}_{1},(\mathcal{B}+\beta_{0}\operatorname{I}_{\mathfrak{a}})\mathcal{P}_{\mathfrak{a}}\mathtt{X}_{2}\rangle.$
2. Note, when $\mathcal{B}=\beta\beta_{0}\operatorname{I}_{\mathfrak{a}}$, singularity of $\operatorname{I}_{\mathfrak{a}}+\beta_{0}^{-1}\mathcal{B}$ means $\beta=-\beta_{0}$, thus, the metric eq. 18 is a multiple of the original bi-invariant metric. In this case, while nominally we deform using $\mathfrak{a}$, there is no actual deformation. For general $\mathcal{B}$, with appropriate conditions, we can likely show if $\operatorname{I}_{\mathfrak{a}}+\beta_{0}^{-1}\mathcal{B}$ is singular, deformation is practically done on a subalgebra $\mathfrak{a}_{1}\subset\mathfrak{a}$. We will not use this unproven statement. 3. The decomposition eq. 20 and the form eq. 21 on $\tilde{\mathfrak{m}}$ are sufficient to define a metric on $\mathcal{G}$ identified with $\tilde{\mathcal{G}}$$\tilde{\mathcal{H}}$, the naturally reductive conditions eq. 8 and eq. 9 can be verified directly without assuming a metric on $\tilde{\mathcal{G}}$. In actuality, we arrive at the decomposition and the metric on $\tilde{\mathfrak{m}}$ assuming $\operatorname{I}_{\mathfrak{a}}+\beta_{0}^{-1}\mathcal{B}$ is nonsingular, following existing literature on the deformation metric [5, 18, 19, 16, 6], then observe that the decomposition and the metric on $\tilde{\mathfrak{m}}$ are both well-defined even when $\operatorname{I}_{\mathfrak{a}}+\beta_{0}^{-1}\mathcal{B}$ is singular. We will present a proof of the theorem along this line, using the bi-invariance metric eq. 22 then show the general case follows from a continuity argument.
@@
From eq. 19, if $(\mathtt{X},\mathtt{A})\in\tilde{\mathfrak{h}}\cap\tilde{\mathfrak{m}}$ then $\mathtt{X}=\mathtt{A}$ and $\hat{\mathcal{B}}\mathtt{A}=\mathtt{A}$, but this shows $\beta_{0}\mathcal{B}^{-1}\mathtt{A}=0$, thus $\mathtt{X}=\mathtt{A}=0$, or $\tilde{\mathfrak{h}}\cap\tilde{\mathfrak{m}}=\{0\}$.
In eq. 20, it is clear we have the equality and that the second line is in $\tilde{\mathfrak{h}}$. To check that the first term is in $\tilde{\mathfrak{m}}$ so that eq. 20 is a direct sum decomposition, we expand
$\hat{\mathcal{B}}(\mathcal{P}_{\mathfrak{a}}\mathtt{X}-(\operatorname{I}_{\mathfrak{a}}+\beta_{0}^{-1}\mathcal{B})\mathcal{P}_{\mathfrak{a}}\mathtt{X}+\beta_{0}^{-1}\mathcal{B}\mathtt{A})=(\mathcal{B}+\beta_{0}\operatorname{I}_{\mathfrak{a}})\mathcal{B}^{-1}(-(\beta_{0}^{-1}\mathcal{B})\mathcal{P}_{\mathfrak{a}}\mathtt{X}+\beta_{0}^{-1}\mathcal{B}\mathtt{A}),$
which simplifies to $(\operatorname{I}_{\mathfrak{a}}+\beta_{0}^{-1}\mathcal{B})\mathcal{P}_{\mathfrak{a}}(\mathtt{X}-\mathtt{A})$, the $\mathcal{A}$-component of the first line.
Applying partial derivative rule on $XA^{-1}$, the differential $d\mathfrak{q}_{(X,A)}$ of $\mathfrak{q}$ applied on direction $(\eta^{\mathcal{G}},\eta^{\mathcal{A}})\in T_{(X,A)}\tilde{\mathcal{G}}$ is given by
(24) $d\mathfrak{q}_{(X,A)}:(\eta^{\mathcal{G}},\eta^{\mathcal{A}})\mapsto X(X^{-1}\eta^{\mathcal{G}}-A^{-1}\eta^{\mathcal{A}})A^{-1}=XA^{-1}\operatorname{Ad}_{A}(X^{-1}\eta^{\mathcal{G}}-A^{-1}\eta^{\mathcal{A}}).$
In particular, at $\operatorname{I}_{\tilde{\mathcal{G}}}$, it maps $(\eta^{\mathcal{G}},\eta^{\mathcal{A}})$ to $\eta^{\mathcal{G}}-\eta^{\mathcal{A}}$. Thus, for $\mathtt{X}\in\mathfrak{g}$, $d\mathfrak{q}_{{}_{\operatorname{I}_{\tilde{\mathcal{G}}}}}:(\mathtt{X},\hat{\mathcal{B}}\mathcal{P}_{\mathfrak{a}}\mathtt{X})\mapsto\mathtt{X}-(\mathcal{B}+\beta_{0}\operatorname{I}_{\mathfrak{a}})\mathcal{B}^{-1}\mathcal{P}_{\mathfrak{a}}\mathtt{X}\in\mathfrak{g}$. Considered as a map from $\tilde{\mathfrak{m}}$ to $\mathfrak{g}$, its inverse applied on $\mathtt{Z}\in\mathfrak{g}$ is the horizontal lift of $\mathtt{Z}$, given by the element $\tilde{\mathtt{Z}}\in\tilde{\mathfrak{m}}$ below
(25) $\tilde{\mathtt{Z}}=(\mathtt{Z}-(\operatorname{I}_{\mathfrak{a}}+\beta_{0}^{-1}\mathcal{B})\mathcal{P}_{\mathfrak{a}}\mathtt{Z},-(\operatorname{I}_{\mathfrak{a}}+\beta_{0}^{-1}\mathcal{B})\mathcal{P}_{\mathfrak{a}}\mathtt{Z}).$
To see this, it is immediate that $d\mathfrak{q}_{{}_{\operatorname{I}_{\tilde{\mathcal{G}}}}}\tilde{\mathtt{Z}}=\mathtt{Z}$. Let $(\mathtt{Z}^{\mathcal{G}},\mathtt{Z}^{\mathcal{A}})$ be the components of $\tilde{\mathtt{Z}}$
(26) $\tilde{\mathtt{Z}}=(\mathtt{Z}^{\mathcal{G}},\mathtt{Z}^{\mathcal{A}})=(\mathtt{Z}+\mathtt{Z}^{\mathcal{A}},\mathtt{Z}^{\mathcal{A}});\quad\mathtt{Z}^{\mathcal{A}}=-(\operatorname{I}_{\mathfrak{a}}+\beta_{0}^{-1}\mathcal{B})\mathcal{P}_{\mathfrak{a}}\mathtt{Z}.$
We show $\tilde{\mathtt{Z}}$ is in $\tilde{\mathfrak{m}}$ by verifying $\hat{\mathcal{B}}\mathcal{P}_{\mathfrak{a}}(\mathtt{Z}+\mathtt{Z}^{\mathcal{A}})=\mathtt{Z}^{\mathcal{A}}$ by simple algebra
$(\mathcal{B}+\beta_{0}\operatorname{I}_{\mathfrak{a}})\mathcal{B}^{-1}\mathcal{P}_{\mathfrak{a}}(\mathtt{Z}-(\operatorname{I}_{\mathfrak{a}}+\beta_{0}^{-1}\mathcal{B})\mathcal{P}_{\mathfrak{a}}\mathtt{Z})=(\mathcal{B}+\beta_{0}\operatorname{I}_{\mathfrak{a}})\mathcal{B}^{-1}(-(\beta_{0}^{-1}\mathcal{B})\mathcal{P}_{\mathfrak{a}}\mathtt{Z})=\mathtt{Z}^{\mathcal{A}}.$
A direct proof of natural reductivity is to expand the $\tilde{\mathfrak{m}}-$form eq. 21 for $\tilde{\mathtt{Z}}$ and verify it simplifies to eq. 18, and similarly verifying $\operatorname{Ad}_{\tilde{\mathcal{H}}}-$invariance and eq. 9 directly. However, as discussed, we give a more comprehensive proof using the metric eq. 22 below.
When $\operatorname{I}_{\mathfrak{a}}+\beta_{0}^{-1}\mathcal{B}$ is nonsingular, the metric eq. 22 is bi-invariant on $\tilde{\mathcal{G}}$, as the $\mathcal{G}$-component is invariant by the bi-invariance of $\mathfrak{g}$, while the $\mathcal{A}$-component is bi-invariant by lemma 2 (for example, it reduces to $(\mathtt{A}_{1},\mathtt{A}_{2})\mapsto\frac{\beta}{\beta+1}\langle\mathtt{A}_{1},\mathtt{A}_{2}\rangle$ if $\mathcal{B}=\beta\beta_{0}\operatorname{I}_{\mathfrak{a}}$). When $\operatorname{I}_{\mathfrak{a}}+\beta_{0}^{-1}\mathcal{B}$ is singular, the metric is not defined on $\tilde{\mathcal{G}}$. However, the form eq. 21 on $\tilde{\mathfrak{m}}$ is still well-defined and continuous in the parameters. Therefore, if the naturally reductive conditions on $\tilde{\mathfrak{m}}$ hold for nonsingular $\operatorname{I}_{\mathfrak{a}}+\beta_{0}^{-1}\mathcal{B}$, the conditions carry to singular values by continuity. Hence, it suffices to prove the nonsingular case, where natural reductivity follows from the bi-invariance of eq. 22, since both conditions eq. 8 and eq. 9 are instances of the bi-invariance condition.
First, we need to show the decomposition eq. 20 is the orthogonal decomposition under eq. 22. The orthogonal complement of $\tilde{\mathfrak{h}}$ under the metric eq. 22 consists of pairs $(\mathtt{X},\mathtt{A})\in\mathfrak{g}\times\mathfrak{a}$ satisfying
$\beta_{0}\langle\mathtt{X},\mathtt{A}_{0}\rangle-\langle\beta_{0}\mathcal{B}(\mathcal{B}+\beta_{0}\operatorname{I}_{\mathfrak{a}})^{-1}\mathtt{A},\mathtt{A}_{0}\rangle=0$
for all $\mathtt{A}_{0}\in\mathfrak{a}$. Since $\langle\mathtt{X},\mathtt{A}_{0}\rangle=\langle\mathcal{P}_{\mathfrak{a}}\mathtt{X},\mathtt{A}_{0}\rangle$, the above reduces to $\mathcal{P}_{\mathfrak{a}}\mathtt{X}-\mathcal{B}(\mathcal{B}+\beta_{0}\operatorname{I}_{\mathfrak{a}})^{-1}\mathtt{A}=0$, or $\mathcal{P}_{\mathfrak{a}}\mathtt{X}=\mathcal{B}(\mathcal{B}+\beta_{0}\operatorname{I}_{\mathfrak{a}})^{-1}\mathtt{A}$, thus $\mathtt{A}=\hat{\mathcal{B}}\mathcal{P}_{\mathfrak{a}}\mathtt{X}$. Hence, the orthogonal complement of $\tilde{\mathfrak{h}}$ is $\tilde{\mathfrak{m}}$, and the metric eq. 22 is nondegenerate on $\tilde{\mathfrak{h}}$ and $\tilde{\mathfrak{m}}$ by lemma 1.
We now verify the metric eq. 21 is the restriction of eq. 22 to $\tilde{\mathfrak{m}}$